OpenStudy (anonymous):

inverse function of x-3/1+4x

OpenStudy (anonymous):

oh my a lot of algebra

OpenStudy (anonymous):

you have to solve \[x=\frac{y-3}{1+4y}\] for \(y\)

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

no i keep getting stuck when i get to x+4xy= y-3 im not sure what to switch

OpenStudy (aum):

\[ x+4xy= y-3 \\ x + 3 + 4xy = y \\ x + 3 = y - 4xy = y(1 - 4x) \\ \frac{x+3}{1-4x} = y \\ y = \frac{x+3}{1-4x} \]

OpenStudy (anonymous):

thanks